# Entering the overcritical regime of nonlinear Breit-Wheeler pair production in collisions of bremsstrahlung $\gamma$-rays and superintense, tightly focused laser pulses

### v1

@@ III.2.2 Total pair yields
Interestingly, the D-model outcomes show optimal values for the focusing with maximum yield at $x\approx 1.5$ for $E_{0}=2.5$ GeV, at $x\approx 2.4$ for $E_{0}=5$ GeV and at $x\approx 3.9$ for $E_{0}=10$ GeV. On the one hand one might have naivly expected that higher $\xi$ and therefore higher intensities would always lead to higher yields, as depicted in Fig. 8 because of the exponential dependency of the rate which rises steeply when $\xi$ grows. But on the other hand the interaction volume decreases for smaller $x$, so that fewer $\gamma$ photons can participate in the creation of pairs; this effect scales quadratically with $x$. For moderate values of $\kappa$, the $x$-dependence of the pair yield is dominated by the depencency of the rate. But for a parameter range where $\kappa$ largely exceeds 1, the rate depency is distinctly flattened and both counteracting effects can compete with each other. Additionally, there is the feature of the D-Model that $\gamma$-photons penetrating the laser pulse decay over time which further reduces the normally beneficial effect of higher intensities on the yield (see Fig. 8 and Fig. 6). This leads to an even greater necessity for a large interaction volume and correspondingly decreased laser intensity, because then the influence of the photon decay is reduced. In summary, maximum focusing is not recommended, but instead it is adviced to aim for specific focusing depending on the incident electron energy to maximise the number of created pairs.
Finally, we draw a comparison with the S-Model. Since this model does not take the $\gamma$-beam decay into account, its predicted balancing of the influence of higher intensity versus larger interaction volume happens at higher degrees of laser focusing than in the D-model. As the dashed red curve in Fig. 10 indicates, the maximum predicted by the S-Model for $E_{0}=2.5$ GeV lies at $x\approx 0.5$. Furthermore the S-Model would yield a maximum at $x\approx 1$ for $E_{0}=5$ GeV and at $x\approx 2$ for $E_{0}=10$ GeV (not shown). By comparison with the corresponding D-model predictions we see that significantly weaker focusing leads to optimum pair yields, when the decay of the $\gamma$-beam is properly taken into account. This is because the decay feature, that leads to a reduction of the pair production relative to the S-model approach, is less pronounced when the field intensity is not too high. A more extended interaction region with somewhat reduced peak intensity is therefore more beneficial.
It is interesting to note that the obtained positions of the focusing optimum in the S-model are proportional to $E_{0}$ and thus correspond to a common maximal value of $\kappa\approx 15$ for all three electron energies. Instead, in the D-model, the location of the maxima grows sublinearly with the increase of $E_{0}$. This behaviour of the D-model—which at first sight might appear surprising, given its characteristic preference for not too high intensities—can be attributed to the mainly contributing range of $\gamma$-photon energies. In Fig. 6 we saw the emergence of a global maximum in the spectrally resolved number of created pairs, which shifts to lower values of $f$, and thus to lower $\kappa$ values when $E_{0}$ increases. This trend may explain why the relative growth of the optimum focusing degree expressed by the parameter $x$ occurs more slowly in the D-model than the S-model. Nevertheless, on an absolute scale, the optimum $x$-values are substantially larger in the D-model than the S-model, corresponding to considerably less tight focusing, as described above.
@@ IV Conclusions
Nonlinear Breit-Wheeler pair production in collisions of a focussed high-intensity laser pulse with a beam of high-energy $\gamma$-photons from bremsstrahlung has been discussed in the overcritical regime $1<\kappa\lesssim 30$, which will be achievable at the next generation laser facilities with intensities of $\sim 10^{23}\ \mathrm{W/cm^{2}}$. The presented results complement our previous study [33] of the intermediate interaction regime around $\kappa\approx 1$. To this end, a theoretical model was used that accounts for the attenuation of the bremsstrahlung beam due to the pair production process. The model coincides with our previous approach in the limit of moderate laser intensities and short pulse durations, but also applies to the overcritical regime.
It was shown that in an experiment combining bremsstrahlung $\gamma$-photons emitted from an incident 10 pC beam of 10 GeV electrons with a 20 fs, $10^{23}$ W/cm2 laser pulse, more than one pair could be generated per shot, which would allow for precision measurements with increased statistical accuracy and deeper understanding of the underlying vacuum structure. That means, if the repetion rate of the laser system is 0.1 Hz, one could produce 360 pairs per hour. The scaling of the pair yield with laser pulse duration was found to be sublinear, because of the $\gamma$-beam decay. Moreover, regarding the impact of bremsstrahlung photons from different parts of the spectrum it was shown that, in contrast to Ref. [33], also $\gamma$-photons of relatively low energy play an important role in the particle creation. Accordingly, by the considered setup a rather broad range of $\kappa$ values is probed simultaneously. Finally, the impact of the laser focusing at fixed laser pulse energy was elucidated, with the outcome that in the overcritical regime there is an optimal focal spot size: Beyond this point, focusing even more tightly to reach higher peak intensity would reduce the pair production yield.
As a general result it has thus been shown that interesting properties of pair creation in the overcritical field regime arise from the relatively slow growth of the production rate with increasing laser intensity: While for small and intermediate values $\kappa\lesssim 1$, the optimum pair yield is known to be achieved for the highest bremsstrahlung energies and the tightest laser focussing, in the considered overcritical regime it is instead more beneficial to relax these conditions in a suitable manner.
@@ Acknowledgements.
This work has been funded by the Deutsche Forschungsgemeinschaft (DFG) under Grant No. 392856280 within the Research Unit FOR 2783/2.

### v2

@@ III.2.2 Total pair yields
Interestingly, the D-model outcomes for the first generation of pairs (dashed lines) show optimal values for the focusing with maximum yield at $x\approx 1.5$ for $E_{0}=2.5$ GeV, at $x\approx 2.4$ for $E_{0}=5$ GeV and at $x\approx 3.9$ for $E_{0}=10$ GeV. On the one hand one might have naivly expected that higher $\xi$ and therefore higher intensities would always lead to higher yields, as depicted in Fig. 8 because of the exponential dependency of the rate which rises steeply when $\xi$ grows. But on the other hand the interaction volume decreases for smaller $x$, so that fewer $\gamma$ photons can participate in the creation of pairs; this effect scales quadratically with $x$. For moderate values of $\kappa$, the $x$-dependence of the pair yield is dominated by the depencency of the rate. But for a parameter range where $\kappa$ largely exceeds 1, the rate depency is distinctly flattened and both counteracting effects can compete with each other. Additionally, there is the feature of the D-model that $\gamma$-photons penetrating the laser pulse decay over time which further reduces the normally beneficial effect of higher intensities on the yield (see Fig. 8 and Fig. 6). This leads to an even greater necessity for a large interaction volume and correspondingly decreased laser intensity, because then the influence of the photon decay is reduced. As a result, the obtained positions of the focusing optimum in the D-model decrease sublinearly with the increase of $E_{0}$.
This feature remains when the contributions from the second pair generation are included (solid lines). These contributions increase for higher electron energies, being almost negligible for 2.5 GeV and most prononunced for 10 GeV, and also increase with tighter focusing, since the values of $\kappa$ are higher then. The relative difference between the black dashed and solid lines amounts to $\approx 20$% at $x\approx 2$, which corresponds to a maximal value of $\kappa\approx 15$ in the bremsstrahlung-laser collision. Although the second generation can enhance the pair yield substantially, it is not large enough to change the overall shape of the curves. Instead the larger interaction volume still plays the crucial role for maximizing the total yield. In result, the inclusion of the second generation only leads for a small shift of the maxima towards a little tighter focusing, with the maximum yield located at $x\approx 1.3$ for 2.5 GeV, $x\approx 2.2$ for 5 GeV and $x\approx 3.3$ for 10 GeV. This implies in summary, that maximum focusing is not recommended, but instead it is adviced to aim for specific focusing depending on the incident electron energy to maximise the number of created pairs.
@@ IV Conclusions
Nonlinear Breit-Wheeler pair production in collisions of a focused high-intensity laser pulse with a beam of high-energy $\gamma$-photons from bremsstrahlung has been discussed in the overcritical regime $1<\kappa\lesssim 30$, which will be achievable at the next generation laser facilities with intensities of $\sim 10^{23}\ \mathrm{W/cm^{2}}$. The presented results complement our previous study [33] of the intermediate interaction regime around $\kappa\approx 1$. To this end, a theoretical model was used that accounts for the attenuation of the bremsstrahlung beam due to the pair production process. This D-model coincides with our previous approach in the limit of moderate laser intensities and short pulse durations, but also applies to the overcritical regime. In particular, the predictions for the total pair yield from both models are in close agreement up to $\kappa\approx 2$, where the CALA experiment [13] is going to operate.
For higher $\kappa$ values the $\gamma$-beam decay has a sizeable effect. One of its consequences is that the first-generation pair yield scales sublinearly with the laser pulse duration. In the considered bremsstrahlung-laser collisions, the D-model was found to represent a good approximation up to $\kappa\approx 6$ for the spectrally resolved pair yields and up to maximal values of $\kappa\approx 12$–15 for the total (bremsstrahlung-integrated) pair yields. However, for larger quantum nonlinearity parameters, the contributions from the second generation of created pairs play a relevant role, which was accounted for within an approximated framework based on Ref. [51]. While the $\gamma$-beam decay tends to lower the pair production, the additional contributions from the second pair generation have an enhancing effect that grows with increasing $\kappa$.
It was shown that in an experiment combining bremsstrahlung $\gamma$-photons emitted from an incident 10 pC beam of 10 GeV electrons with a 20 fs, $10^{23}$ W/cm2 laser pulse, about 1.7 pairs could be generated per shot (to which the second pair generation gives a relative contribution of approximately 30%). Hence, if the repetion rate of the laser system is 0.1 Hz, about 600 pairs could be produced per hour. This would allow for precision measurements with increased statistical accuracy and deeper understanding of the underlying vacuum structure.
Regarding the impact of bremsstrahlung photons from different parts of the spectrum it was shown that, in contrast to Ref. [33], also $\gamma$-photons of relatively low energy play an important role in the particle creation. Accordingly, by the considered setup a rather broad range of $\kappa$ values can be probed simultaneously. Finally, the impact of the laser focusing at fixed laser pulse energy was elucidated, with the outcome that in the overcritical regime there is an optimal focal spot size: Beyond this point, focusing even more tightly to reach higher peak intensity would reduce the pair production yield.
@@ Acknowledgements.
This work has been funded by the Deutsche Forschungsgemeinschaft (DFG) under Grant No. 392856280 within the Research Unit FOR 2783/2. Some of the numerical computations were carried out by using the code [52] for photon-seeded pair production of M. Pouyez.